Calculate m if 800g solid rises by 4oC when 4000J is supplied
1
Expert's answer
2021-08-29T16:56:09-0400

I supposed m means specific heat capacity (let's denote is "c"). By definition, it is given as follows:


"c = \dfrac{Q}{m\Delta T}"

where "Q = 4000J" is the heat supplied, "m = 800g = 0.8kg" is the mass of the soil, and "\Delta T = 4\degree C" is the rise in the temperature. Thus, obtain:


"c = \dfrac{4000J}{0.8kg\cdot 4\degree C} = 1250\dfrac{J}{kg\cdot \degree C}"

Answer. "1250\dfrac{J}{kg\cdot \degree C}" .